In a sample of 300 students, 68% said they own an iPod and a smart phone. Compute a 95% confidence interval for the true percent of students who own an iPod and a smartphone.


1
Expert's answer
2021-06-07T03:41:52-0400

"p=0.68 \\

q=1-p=1-0.68=0.32 \\

CI \; 0.95 \\

α=1-0.95=0.05 \\

\frac{α}{2}=0.025 \\

Z_{0.025}=1.959 \\

CI = p ±Z_{0.025} \sqrt{\frac{pq}{n}} \\

= 0.68 ± 1.959 \times \sqrt{\frac{0.68 \times 0.32}{300}} \\

= 0.68 ±0.052"

CI: 0.628 ≤ p ≤0.732